Cadbury Company uses a job order costing system. Manufacturing overhead is applied on the basis of direct labor cost. Total manufacturing overhead was estimated to be $144,020 for the year; direct labor was estimated to total $151,600.

(1/1) (12/31)
Raw Materials Inventory $ 14,400 $ 9,900
Work in Process Inventory $ 19,300 $ 22,100
Finished Goods Inventory $ 42,400 $ 30,600


The following transactions have occurred during the year.

Raw materials purchases $ 110,000
Direct materials used $ 103,100
Direct labor $ 127,500
Indirect materials used $ 11,400
Indirect labor $ 16,100
Factory equipment depreciation $ 26,600
Factory rent $ 21,300
Factory utilities $ 8,600
Other factory costs $ 7,600

a) Calculate the predetermined overhead rate

Predetermined Overhead Rate    %

(b) Calculate cost of goods manufactured.

Cost of Goods Manufactured

c) Calculate the over- or underapplied overhead. (Input the amount as positive value.)

  


(d) Calculate adjusted cost of goods sold

Adjusted Cost of Goods Sold   
Answer #1

A.Predetermined overhead rate = total manufacturing overhead/ direct labor hours

144020/151600=

95%
B
Cost of Goods manufactured
Direct materials:
Material used in production 103100
Direct labor 127500
Manufacturing overhead applied( on direct labor(127500*95%)) 121125
Add: Beginning Work in process 19300
Less: Ending Work in process 22100
Cost of goods manufactured 348925
C

Overhead Underapplied =

(Actual manufacturing overhead - Estimated manufacturing overhead) refer Working note at End.

(11400+16100+26600+21300+8600+7600)=91600-121125

29525
D
Cost of goods manufactured 348925
Add: Beginning Finished inventory 42400
Cost of goods available for sale 391325
Less: Ending finished inventory 30600
Cost of Goods Sold 360725
Add: Underapplied overhead 29525
Adjusted Cost of Goods Sold 390250
Working note:
Total actual overhead
Indirect materials used $ 11,400
Indirect labor $ 16,100
Factory equipment depreciation $ 26,600
Factory rent $ 21,300
Factory utilities $ 8,600
Other factory costs $ 7,600
Total 91600